# Differentiable Structure Learning for General Binary Data

### v1

@@ 4.3 Minimal equivalence class
For the procedure in Section 4.2, if we retains those pairs $(f_{\pi},G_{\pi})$ with the fewest edges, thereby recovering ${\mathcal{E}}_{\mathrm{min}}(\bm{p})$. Minimality is closely related to the Sparest Markov Representation (SMR) assumption (Raskutti and Uhler, 2018; Lam et al., 2022), which posits that, among all DAGs compatible with a distribution, the sparsest one is unique up to Markov equivalence, i.e., encoding the same conditional independence relationship. SMR is weaker than the well-known faithfulness assumption (Peters et al., 2017): if a distribution $P$ is faithful to a DAG $G$, then $(G,P)$ satisfies SMR. Hence, targeting the sparsest graph is both principled and practically appealing. Let ${\mathcal{M}}(G)$ denote the Markov Equivalence class of a DAG $G$, i.e. the set of all DAGs encoding the same conditional‐independence relations. Formal definitions of these concepts appears in Appendix C.6.
@@ 5 Continuous structure learning for general binary data
$\begin{split}\mathbf{H}_{j}=&(\underbrace{h^{j,0}}_{\text{constant}},\underbrace{h^{j,1},\ldots,h^{j,p}}_{\text{first order }},\underbrace{h^{j,12},\ldots,h^{j,(p-1)p}}_{\text{second order}},\underbrace{\ldots}_{\text{third to }(p-1)\text{-th order}},\underbrace{h^{j,123\ldots p}}_{p\text{-th order}})^{\top}\in\mathbb{R}^{2^{p}}\end{split}$ (9)
@@
$[W(\mathbf{H})]_{ij}=\sum_{S\subseteq[p],i\in S}\left(h^{j,S}\right)^{2}$ (10)
so $W(\mathbf{H})]_{ij}>0$ if and only if some interaction involving $X_{i}$ contributes to the equation for $X_{j}$, that is some coefficient $h^{j,S}$ with $i\in S$ is non–zero. Self-loops are forbidden, so we impose
$h^{j,S}=0\quad\text{whenever }j\in S,\qquad\forall\,j\in[p],\;S\subseteq[p].$
@@
where $\mathbbm{1}\{\cdot\}$ is the indicator function. $p_{\lambda,\delta}$ is quadratic on $0,\delta]$ and flat beyond $\delta$, smoothly approximating an $\ell_{0}$ penalty and thereby encouraging a sparse $W(\mathbf{H})$ without sacrificing differentiability. Let $p_{\lambda,\delta}(W(\mathbf{H}))=\sum_{i\neq j}p_{\lambda,\delta}([W(\mathbf{H})]_{ij})$. Define the regularized score
@@
$\begin{split}\min_{\mathbf{H}}\qquad&s(\mathbf{H};\lambda,\delta,{\mathbf{X}})\\ \text{subject to }\qquad&h(W(\mathbf{H}))=0\\ \qquad&h^{j,S}=0\text{ if }j\in S\quad\forall j\in[p],\forall S\subseteq[p]\end{split}$ (13)
@@
For any $\bm{p}>0$, there exist $\lambda,\delta>0$ sufficiently small such that ${\mathcal{O}}_{\infty,\lambda,\delta}={\mathcal{E}}_{\min}(\bm{p}).$ Moreover, under Sparest Markov Representation (or faithfulness) assumption, for any $(\bm{f}_{\pi_{1}},G_{\pi_{1}}),(\bm{f}_{\pi_{2}},G_{\pi_{2}})\in{\mathcal{O}}_{\infty,\lambda,\delta}$, it holds ${\mathcal{M}}(G_{\pi_{1}})={\mathcal{M}}(G_{\pi_{2}})$.
@@ 6 Experiments
We solve (13) using the DAGMA optimization framework of Bello et al. (2022), which tackles a sequence of unconstrained problems by incorporating the acyclicity constraint as a penalty term with an increasing weight. Our method—denoted dagma-ho (for “higher‐order”)—is compared against several baselines, including DAGMA (Bello et al., 2022), PC (Spirtes and Glymour, 1991), and FGES (Ramsey et al., 2017). Our primary empirical results are shown in Figure 1 and 2. We use the structural Hamming distance (SHD) as the main metric to evaluate the difference between the estimated and the ground truth graph. Lower SHD indicates better estimation accuracy. Given the multivariate Bernoulli distribution is nonidentifiable, we compare the completed partially directed acyclic graph (CPDAG) of the estimated and ground truth graph. Full experimental details are provided in Appendix D.
@@
In Figure 1(a), we simulate data $\mathbf{X}$} according to the SEM in (7), including both first‐order effects and second‐order interactions among parents whenever a node has multiple parents. We observe that dagma-ho (ours) achieves competitive performance relative to the baselines. In contrast, dagma-1st (the original DAGMA), which models only first‐order effects, exhibits markedly poor recovery when second‐order interactions are present. These results underscore that a purely linear form is insufficient to capture the complex dependencies in general discrete data and can lead to substantial estimation errors whenever higher‐order terms play a role. In Figure 1(b), we repeat the simulation while ablating all first‐order effects for nodes with multiple parents, retaining only the highest‐order interaction term. Under these conditions, dagma-ho (ours) consistently attains near‐optimal recovery performance across all settings, showing the robustness of our method.
@@
In Figure 2, we consider a larger DAG, where including all possible interaction terms would cause the feature map $\Phi(\mathbf{X})$ to grow exponentially and render the optimization in (13) intractable. To address this, we propose a two-stage approach, which we call Notears-MLP-reg. First, we adapt the original NOTEARS-MLP framework of Zheng et al. (2020) to handle discrete data, and use it to learn an initial graph. From this graph we extract a topological ordering $\pi$. In the second (finetuning) stage, we follow the procedure of Section 4.2: we construct only those interaction features consistent with $\pi$, and then fit logistic regressions with first order and second order to recover the final edge structure $G$. Notears-MLP-reg attains performance on par with existing baselines. While all methods degrade as graph size and density increase, our approach exhibits greater robustness and consistently recovers meaningful structures.
Further details on the experimental setup and additional results can be found in the appendix.
@@
Figure 1: Results in terms of SHD between MECs of estimated graph and ground truth. Lower is better. Column: $k=\{1,2\}$. Row: random graph types. $\{\text{ER,SF}\}$-$k=\{\text{Scale-Free},\text{Erdős–Rényi}\}$ graphs with kdkd expected edges. Here $p=\{5,6,7,8,9\}$. DAGMA (Bello et al., 2022) is renamed as “DAGMA-1st”, to emphasize only linear term is used. Error bars denote the standard error computed over 10 replications.
Figure 2: Results in terms of SHD between MECs of estimated graph and ground truth. Lower is better. Column: $k=\{1,2,4\}$. Row: random graph types. $\{\text{ER,SF}\}$-$k=\{\text{Scale-Free},\text{Erdős–Rényi}\}$ graphs with kdkd expected edges. Here $p=\{10,20,30,40\}$. Notears-mlp-reg is our two stage approach. DAGMA (Bello et al., 2022) is renamed as “DAGMA-1st”, to emphasize only linear term is used.
@@ 7 Conclusion
To scale beyond small graphs, we further propose a practical two-stage heuristic, Notears-MLP-reg, which first learns a coarse structure via NOTEARS-MLP adapted to discrete data and then refines edge estimates by fitting logistic regressions along the inferred topological order. While this approach yields strong numerical performance on larger networks, its theoretical properties remain unexplored. In future work, we aim to develop principled, scalable algorithms for discrete structure learning that retain both computational tractability and rigorous identifiability guarantees in regimes with thousands of nodes.
@@ Appendix A Preliminary Technical Results
$X_{S}\sim\mathrm{MultiBernoulli}(\bm{p}_{S}),$
where $\bm{p}_{S}>0$.
@@
$\bm{p}>0\quad\Longleftrightarrow\quad|\bm{f}|<\infty.$
@@
$\ell(w)\;=\;\mathbb{E}\!\Bigl[\log\bigl(1+\exp\bigl(w^{\top}\Phi(X_{\pi(1)},\dots,X_{\pi(j-1)})\bigr)\bigr)\;-\;X_{\pi(j)}\,w^{\top}\Phi(X_{\pi(1)},\dots,X_{\pi(j-1)})\Bigr],\quad w\in\mathbb{R}^{2^{\,j-1}}.$
Then $\ell(w)$ is strictly convex and therefore admits a unique minimizer $w^{*}_{\pi,j}$.
@@
$L_{c}=\left\{x\in\mathbb{R}^{p}\mathrel{\mathop{\ordinarycolon}}f(x)\leq c\right\}$
@@ B.2 Proof of Lemma 2
$\begin{split}&\text{p($j_{1},j_{2},\ldots,j_{r}$ positions are one, others are zero)}\\ =&\frac{\exp(S^{j_{1}j_{2}\cdots j_{r}})}{\exp(b(\bm{f}))}.\\ =&\frac{\exp(S^{j_{1}j_{2}\cdots j_{r}})}{\sum_{r=1}^{K}\left[1+\left(\sum_{1\leq j_{1}<j_{2}<\ldots<j_{r}\leq K}\exp[S^{j_{1}j_{2}\ldots j_{r}}]\right)\right]}\end{split}$ (22)
@@ B.3 Proof of Lemma 3
The marginal distribution of $(X_{\pi(1)},\ldots,X_{\pi(j-1)})$ is still a multivariate Bernoulli distribution, with natural parameters $\bm{p}_{\pi,j-1}$. i.e., $(X_{\pi(1)},\ldots,X_{\pi(j-1)})\sim MultiBernoulli(\bm{p}_{\pi,j-1})$. By lemma 1, $q_{\min,j-1}=\min_{i\in[2^{j-1}]}[\bm{p}_{\pi,j-1}]_{i}>0$. Hence for any nonzero $\beta\in\mathbb{R}^{2^{\,j-1}}$,
@@
because the sum is over all $z\in\{0,1\}^{2^{j-1}}$. So $(\beta^{\top}z)^{2}>0$ for at least one $z$ when $\beta\neq 0$.
Since the Hessian is everywhere positive definite, $\ell(w)$ is strictly convex and therefore has a unique minimizer.
@@ B.6 Proof of Theorem 3
$[W(\mathbf{H})]_{ij}$ $=\sum_{S\subseteq[p],\,i\in S}\bigl(h^{j,S}\bigr)^{2}.$ (39)
@@
$[W(\mathbf{H})]_{ij}$ $=\sum_{S\subseteq[p],\,i\in S}\bigl|h^{j,S}\bigr|.$ (40)
Let $\mathbf{H}_{1}$ and $\mathbf{H}_{2}$ be two parameter values. We show there exists $L$ such that
@@
$\|W(\mathbf{H}_{1})-W(\mathbf{H}_{2})\|_{2}$ $=\sqrt{\sum_{j}\sum_{i}\Bigl(\sum_{S\subseteq[p],\,i\in S}\bigl|h_{1}^{j,S}\bigr|-\sum_{S\subseteq[p],\,i\in S}\bigl|h_{2}^{j,S}\bigr|\Bigr)^{2}}\,.$ (42)
@@
$\|\mathbf{H}_{1}-\mathbf{H}_{2}\|_{2}$ $=\sqrt{\sum_{j}\sum_{S\subseteq[p]}\bigl(h_{1}^{j,S}-h_{2}^{j,S}\bigr)^{2}}\,.$ (43)
@@
$\begin{split}\left(\sum_{S\subseteq[p],i\in S}\mathinner{\!\left\lvert h_{1}^{j,S}\right\rvert}-\sum_{S\subseteq[p],i\in S}\mathinner{\!\left\lvert h_{2}^{j,S}\right\rvert}\right)^{2}=&\left(\sum_{S\subseteq[p],i\in S}\left(\mathinner{\!\left\lvert h_{1}^{j,S}\right\rvert}-\mathinner{\!\left\lvert h_{2}^{j,S}\right\rvert}\right)\right)^{2}\\ \leq&|S\subseteq[p],i\in S|\sum_{S\subseteq[p],i\in S}\left(\mathinner{\!\left\lvert h_{1}^{j,S}\right\rvert}-\mathinner{\!\left\lvert h_{2}^{j,S}\right\rvert}\right)^{2}\\ \leq&2^{p-1}\sum_{S\subseteq[p],i\in S}\left({h_{1}^{j,S}}-{h_{2}^{j,S}}\right)^{2}\\ \leq&2^{p-1}\sum_{S\subseteq[p]}\left({h_{1}^{j,S}}-{h_{2}^{j,S}}\right)^{2}\end{split}$ (44)
@@
$\begin{split}\|W(\mathbf{H}_{1})-W(\mathbf{H}_{2})\|_{2}=&\sqrt{\sum_{j}\sum_{i}\left(\sum_{S\subseteq[p],i\in S}\mathinner{\!\left\lvert h_{1}^{j,S}\right\rvert}-\sum_{S\subseteq[p],i\in S}\mathinner{\!\left\lvert h_{2}^{j,S}\right\rvert}\right)^{2}}\\ \leq&\sqrt{\sum_{j}\sum_{i}2^{p-1}\sum_{S\subseteq[p]}\left({h_{1}^{j,S}}-{h_{2}^{j,S}}\right)^{2}}\\ \leq&\sqrt{2^{p-1}p\sum_{j}\sum_{S\subseteq[p]}\left({h_{1}^{j,S}}-{h_{2}^{j,S}}\right)^{2}}\\ =&\sqrt{p}2^{(p-1)/2}\sqrt{\sum_{j}\sum_{S\subseteq[p]}\left(h_{1}^{j,S}-h_{2}^{j,S}\right)^{2}}\\ =&\sqrt{p}2^{(p-1)/2}\|\mathbf{H}_{1}-\mathbf{H}_{2}\|_{2}\end{split}$ (45)

### v2

@@ 4.3 Minimal equivalence class
For the procedure in Section 4.2, if we retains those pairs $(f_{\pi},G_{\pi})$ with the fewest edges, thereby recovering ${\mathcal{E}}_{\mathrm{min}}(\bm{p})$. Minimality is closely related to the Sparsest Markov Representation (SMR) assumption (Raskutti and Uhler, 2018; Lam et al., 2022), which posits that, among all DAGs compatible with a distribution, the sparsest one is unique up to Markov equivalence, i.e., encoding the same conditional independence relationship. SMR is weaker than the well-known faithfulness assumption (Peters et al., 2017): if a distribution $P$ is faithful to a DAG $G$, then $(G,P)$ satisfies SMR. Hence, targeting the sparsest graph is both principled and practically appealing. Let ${\mathcal{M}}(G)$ denote the Markov Equivalence class of a DAG $G$, i.e. the set of all DAGs encoding the same conditional‐independence relations. Formal definitions of these concepts appears in Appendix C.6.
@@ 5 Continuous structure learning for general binary data
$\begin{split}\mathbf{H}_{j}=&(\underbrace{h^{0,j}}_{\text{constant}},\underbrace{h^{1,j},\ldots,h^{p,j}}_{\text{first order }},\underbrace{h^{12,j},\ldots,h^{(p-1)p,j}}_{\text{second order}},\underbrace{\ldots}_{\text{third to }(p-1)\text{-th order}},\underbrace{h^{123\ldots p,j}}_{p\text{-th order}})^{\top}\in\mathbb{R}^{2^{p}}\end{split}$ (9)
@@
$[W(\mathbf{H})]_{ij}=\sum_{S\subseteq[p],i\in S}\left(h^{S,j}\right)^{2}$ (10)
so $W(\mathbf{H})]_{ij}>0$ if and only if some interaction involving $X_{i}$ contributes to the equation for $X_{j}$, that is some coefficient $h^{S,j}$ with $i\in S$ is non–zero. Self-loops are forbidden, so we impose
$h^{S,j}=0\quad\text{whenever }j\in S,\qquad\forall\,j\in[p],\;S\subseteq[p].$
@@
where $\mathbbm{1}(\cdot)$ is the indicator function. $p_{\lambda,\delta}$ is quadratic on $0,\delta]$ and flat beyond $\delta$, smoothly approximating an $\ell_{0}$ penalty and thereby encouraging a sparse $W(\mathbf{H})$ without sacrificing differentiability. Let $p_{\lambda,\delta}(W(\mathbf{H}))=\sum_{i\neq j}p_{\lambda,\delta}([W(\mathbf{H})]_{ij})$. Define the regularized score
@@
$\begin{split}\min_{\mathbf{H}}\qquad&s(\mathbf{H};\lambda,\delta,{\mathbf{X}})\\ \text{subject to }\qquad&h(W(\mathbf{H}))=0\\ \qquad&h^{S,j}=0\text{ if }j\in S\quad\forall j\in[p],\forall S\subseteq[p]\end{split}$ (13)
@@
For any $\bm{p}>0$, there exist $\lambda,\delta>0$ sufficiently small such that ${\mathcal{O}}_{\infty,\lambda,\delta}={\mathcal{E}}_{\min}(\bm{p}).$ Moreover, under Sparsest Markov Representation (or faithfulness) assumption, for any $(\bm{f}_{\pi_{1}},G_{\pi_{1}}),(\bm{f}_{\pi_{2}},G_{\pi_{2}})\in{\mathcal{O}}_{\infty,\lambda,\delta}$, it holds ${\mathcal{M}}(G_{\pi_{1}})={\mathcal{M}}(G_{\pi_{2}})$.
@@ 6 Experiments
We solve (13) using either DAGMA (Bello et al., 2022) or NOTEARS-MLP (Zheng et al., 2020). Both methods tackle a sequence of unconstrained problems by incorporating the acyclicity constraint as a penalty term with an increasing weight. [^3] Concretely, we (i) augment DAGMA with higher-order terms on small graphs to pursue exact recovery, and (ii) employ NOTEARS-MLP to obtain a topological order on large graphs and subsequently fit logistic regressions for edge selection. We use these two pathways because, in our experiments, they strike different speed–accuracy trade-offs. See Appendix D.2 for implementation details. Our method—denoted BiNOTEARS—is compared against several baselines, including DAGMA (Bello et al., 2022), PC (Spirtes and Glymour, 1991), and FGES (Ramsey et al., 2017). Our primary empirical results appear in Figures 1 and 2. We evaluate accuracy using structural Hamming distance (SHD), which measures the discrepancy between the estimated and ground-truth graphs, with lower values indicating better performance. Because the multivariate Bernoulli model is nonidentifiable, we compare completed partially directed acyclic graphs (CPDAGs) of the estimates and the ground truth. Full experimental details are provided in Appendix D.
@@
In Figure 1 (a), we simulate data $\mathbf{X}$} according to the SEM in (7) for small DAG ($d<10$), including both first‐order effects and second‐order interactions among parents whenever a node has multiple parents. We observe that BiNOTEARS (ours) achieves competitive performance relative to the baselines. In contrast, DAGMA which models only first‐order effects, exhibits markedly poor recovery when second‐order interactions are present. These results underscore that a purely linear form is insufficient to capture the complex dependencies in general discrete data and can lead to substantial estimation errors whenever higher‐order terms play a role. In Figure 1 (b), we repeat the simulation while ablating all first‐order effects for nodes with multiple parents, retaining only the highest‐order interaction term. Under these conditions, BiNOTEARS (ours) consistently attains near‐optimal recovery performance across all settings, showing the robustness of our method.
@@
In Figure 2, we consider a larger DAG, where including all possible interaction terms would cause the feature map $\Phi(\mathbf{X})$ to grow exponentially and render the optimization in (13) intractable. To address this, our BiNOTEARS works in a two-stage fashion. First, we adapt the original NOTEARS-MLP framework of Zheng et al. (2020) to handle discrete data, and use it to learn an initial graph. From this graph we extract a topological ordering $\pi$. In the second (finetuning) stage, we follow the procedure of Section 4.2: we construct only those interaction features consistent with $\pi$, and then fit logistic regressions with first order and second order to recover the final edge structure $G$. BiNOTEARS attains performance on par with existing baselines. While all methods degrade as graph size and density increase, our approach exhibits robustness and consistently recovers meaningful structures.
We further evaluate NOTEARS (linear), NOTEARS-MLP, and BiNOTEARS (ours) on the real-world dataset of Sachs et al. (2005). The data contain n=7,466n=7{,}466 measurements of $d=11$ proteins and phospholipids in human immune cells and are accompanied by a widely used consensus network that serves as a gold standard. Results are summarized in the Table. Further details on the experimental setup and additional results can be found in the Appendices D and E.
Method SHD Num. of Edges
NOTEARS (Zheng et al., 2018) 22 18
NOTEARS-MLP (Zheng et al., 2020) 16 13
BiNOTEARS (Ours) 13 15
@@
Figure 1: Results in terms of SHD between MECs of estimated graph and ground truth. Lower is better. Column: $k=\{1,2\}$. Row: random graph types. $\{\text{ER,SF}\}$-$k=\{\text{Scale-Free},\text{Erdős–Rényi}\}$ graphs with $k\cdot d$ expected edges. Here $p=\{5,6,7,8,9\}$. Error bars denote the standard error computed over 10 replications.
Figure 2: Results in terms of SHD between MECs of estimated graph and ground truth. Lower is better. Column: $k=\{1,2,4\}$. Row: random graph types. $\{\text{ER,SF}\}$-$k=\{\text{Scale-Free},\text{Erdős–Rényi}\}$ graphs with kdkd expected edges. Here $p=\{10,20,30,40\}$. BiNOTEARS is our two-stage approach.
@@ 7 Conclusion
To scale beyond small graphs, we further propose a practical two-stage heuristic, BiNOTEARS, which first learns a coarse structure via NOTEARS-MLP adapted to discrete data and then refines edge estimates by fitting logistic regressions along the inferred topological order. While this approach yields strong numerical performance on larger networks, its theoretical properties remain unexplored. In future work, we aim to develop principled, scalable algorithms for discrete structure learning that retain both computational tractability and rigorous identifiability guarantees in regimes with thousands of nodes.
@@ Appendix A Preliminary Technical Results
$X_{S}\sim\mathrm{MultiBernoulli}(\bm{p}_{S}),$ (16)
for certain $\bm{p}_{S}\in\mathbb{R}^{|S|}$ where $\bm{p}_{S}>0$.
@@
$\bm{p}>0\quad\Longleftrightarrow\quad|\bm{f}|<\infty.$ (17)
@@
$\ell(w)=\mathbb{E}\!\Bigl[\log\bigl(1+\exp\bigl(w^{\top}\Phi(X_{\pi(1)},\dots,X_{\pi(j-1)})\bigr)\bigr)-X_{\pi(j)}\,w^{\top}\Phi(X_{\pi(1)},\dots,X_{\pi(j-1)})\Bigr]$ (18)
where $w\in\mathbb{R}^{2^{\,j-1}}$. Then $\ell(w)$ is strictly convex and therefore admits a unique minimizer $w^{*}_{\pi,j}$.
@@
$L_{c}=\left\{x\in\mathbb{R}^{p}\mathrel{\mathop{\ordinarycolon}}f(x)\leq c\right\}$ (20)
@@ B.2 Proof of Lemma 2
$\begin{split}&P\text{($j_{1},j_{2},\ldots,j_{r}$ positions are one, others are zero)}\\ =&\frac{\exp(S^{j_{1}j_{2}\cdots j_{r}})}{\exp(b(\bm{f}))}.\\ =&\frac{\exp(S^{j_{1}j_{2}\cdots j_{r}})}{\sum_{r=1}^{K}\left[1+\left(\sum_{1\leq j_{1}<j_{2}<\ldots<j_{r}\leq K}\exp[S^{j_{1}j_{2}\ldots j_{r}}]\right)\right]}\end{split}$ (27)
@@ B.3 Proof of Lemma 3
The marginal distribution of $(X_{\pi(1)},\ldots,X_{\pi(j-1)})$ is still a multivariate Bernoulli distribution, with natural parameters $\bm{p}_{\pi,j-1}$, i.e., $(X_{\pi(1)},\ldots,X_{\pi(j-1)})\sim MultiBernoulli(\bm{p}_{\pi,j-1})$. By lemma 1, $q_{\min,j-1}=\min_{i\in[2^{j-1}]}[\bm{p}_{\pi,j-1}]_{i}>0$. Hence for any nonzero $\beta\in\mathbb{R}^{2^{\,j-1}}$,
@@
Because the sum is over all $z\in\{0,1\}^{2^{j-1}}$. So $(\beta^{\top}z)^{2}>0$ for at least one $z$ when $\beta\neq 0$. Since the Hessian is everywhere positive definite, $\ell(w)$ is strictly convex and therefore has a unique minimizer.
@@ B.6 Proof of Theorem 3
$[W(\mathbf{H})]_{ij}$ $=\sum_{S\subseteq[p],\,i\in S}\bigl(h^{S,j}\bigr)^{2}.$ (44)
@@
$[W(\mathbf{H})]_{ij}$ $=\sum_{S\subseteq[p],\,i\in S}\bigl|h^{S,j}\bigr|.$ (45)
The reason of equivalence is that they encode the same nonzero pattern of $W(\mathbf{H})]_{ij}$. Let $\mathbf{H}_{1}$ and $\mathbf{H}_{2}$ be two parameter values. We show there exists $L$ such that
@@
$\|W(\mathbf{H}_{1})-W(\mathbf{H}_{2})\|_{2}$ $=\sqrt{\sum_{j}\sum_{i}\Bigl(\sum_{S\subseteq[p],\,i\in S}\bigl|h_{1}^{S,j}\bigr|-\sum_{S\subseteq[p],\,i\in S}\bigl|h_{2}^{S,j}\bigr|\Bigr)^{2}}\,.$ (47)
@@
$\|\mathbf{H}_{1}-\mathbf{H}_{2}\|_{2}$ $=\sqrt{\sum_{j}\sum_{S\subseteq[p]}\bigl(h_{1}^{S,j}-h_{2}^{S,j}\bigr)^{2}}\,.$ (48)
@@
$\begin{split}\left(\sum_{S\subseteq[p],i\in S}\mathinner{\!\left\lvert h_{1}^{S,j}\right\rvert}-\sum_{S\subseteq[p],i\in S}\mathinner{\!\left\lvert h_{2}^{S,j}\right\rvert}\right)^{2}=&\left(\sum_{S\subseteq[p],i\in S}\left(\mathinner{\!\left\lvert h_{1}^{S,j}\right\rvert}-\mathinner{\!\left\lvert h_{2}^{S,j}\right\rvert}\right)\right)^{2}\\ \leq&|S\subseteq[p],i\in S|\sum_{S\subseteq[p],i\in S}\left(\mathinner{\!\left\lvert h_{1}^{S,j}\right\rvert}-\mathinner{\!\left\lvert h_{2}^{S,j}\right\rvert}\right)^{2}\\ \leq&2^{p-1}\sum_{S\subseteq[p],i\in S}\left({h_{1}^{S,j}}-{h_{2}^{S,j}}\right)^{2}\\ \leq&2^{p-1}\sum_{S\subseteq[p]}\left({h_{1}^{S,j}}-{h_{2}^{S,j}}\right)^{2}\end{split}$ (49)
@@
$\begin{split}\|W(\mathbf{H}_{1})-W(\mathbf{H}_{2})\|_{2}=&\sqrt{\sum_{j}\sum_{i}\left(\sum_{S\subseteq[p],i\in S}\mathinner{\!\left\lvert h_{1}^{S,j}\right\rvert}-\sum_{S\subseteq[p],i\in S}\mathinner{\!\left\lvert h_{2}^{S,j}\right\rvert}\right)^{2}}\\ \leq&\sqrt{\sum_{j}\sum_{i}2^{p-1}\sum_{S\subseteq[p]}\left({h_{1}^{S,j}}-{h_{2}^{S,j}}\right)^{2}}\\ \leq&\sqrt{2^{p-1}p\sum_{j}\sum_{S\subseteq[p]}\left({h_{1}^{S,j}}-{h_{2}^{S,j}}\right)^{2}}\\ =&\sqrt{p}2^{(p-1)/2}\sqrt{\sum_{j}\sum_{S\subseteq[p]}\left(h_{1}^{S,j}-h_{2}^{S,j}\right)^{2}}\\ =&\sqrt{p}2^{(p-1)/2}\|\mathbf{H}_{1}-\mathbf{H}_{2}\|_{2}\end{split}$ (50)